OpenStudy (anonymous):

Which of the following can be expressed as the sum of the 2 composites? A. 5 B. 15 C. 28

OpenStudy (calculusxy):

Do you mean the two composites out of the three?

OpenStudy (calculusxy):

If so then it will be B and C.

OpenStudy (anonymous):

The sum of two composites, it can be more than one choice.

Directrix (directrix):

I am not sure if this "the 2 composites" means any two composites or two particular composites. 15 = 6 + 9 28 = 24 + 4

Directrix (directrix):

@wwhitney94

OpenStudy (calculusxy):

I have seen this way when summing two composites.